# General properties of the electric Penrose process

### v1

@@ VIII.1 Case (i), turning point for particle 2 (TP2)
VIII.2 Case (i), turning point for particle 0 (TP0)
@@
Let particles 1 and 2 be ejected in point $r_{d}$ along the trajectory of particle 0. Now, it follows from (45) that either $L_{0}=0$ or $P_{1}=0=P_{2}$. In the first case, $\tilde{m}_{0}=m_{0}$, so
@@
In the second case the point $r_{d}$ is the turning point for all three particles and we turn to this possibility in the next subsection.
VIII.3 Case (ii), turning point for all three particles (TP3)
@@
$E_{2}=q_{2}\varphi+\frac{b_{2}}{2m_{0}^{2}}\tilde{m}_{0}N+\frac{\sqrt{d}\left|L_{0}\right|}{2m_{0}^{2}r_{d}}\text{,}$ (108)
$E_{1}=q_{1}\varphi+\frac{b_{1}}{2m_{0}^{2}}\tilde{m}_{0}N-\frac{\sqrt{d}\left|L_{0}\right|}{2m_{0}^{2}r_{d}}\text{.}$ (109)
@@
$L_{2}=\frac{1}{2m_{0}^{2}}(L_{0}b_{2}+signL_{0}\sqrt{d}\tilde{m}_{0}r_{d})=L_{2}^{(+)},$ (110)
$L_{1}=\frac{1}{2m_{0}^{2}}(L_{0}b_{1}-signL_{0}\sqrt{d}\tilde{m}_{0}r_{d})=L_{1}^{(-)}\text{,}$ (111)
@@
IX Threshold for the Penrose process
@@
X Velocities and gamma factors for scenario I
@@
$\gamma_{01}=\frac{b_{1}}{2m_{0}m_{1}}\text{, }v_{1}=\frac{\sqrt{d}}{b_{1}}\text{.}$ (123)
@@
XI Scenario B and the maximum of efficiency
@@
Here, as before, particle 2 moves to infinity (now in the same direction as particle 0) and particle 1 moves in the inward direction (opposite to particle 0). However, now the signs before the radicals are opposite to those in (127), (128). The condition for the PP now gives us
@@
XII Relation to the BSW effect
@@
XIII Summary and Conclusions
@@
We formulated the results in two forms. The first one includes dependence of the outcome of masses of particles. The second form expresses them in terms of velocities and gamma factors of relative motion between a parent particle and daughter ones. The corresponding formulas obtained for a static charged metric are similar to whose obtained in wald for rotating metrics and neutral particles. Meanwhile, there are some essential differences. For the Penrose process to occur in the Kerr or other rotating metric, there is a lower bound on velocity of a daughter particle in the center of mass frame, this bound being rather high wald . This creates obstacles for using this process. Meanwhile, in our case, the situation is more diverse and there are scenarios in which the velocity obeys the upper (not lower) bound, so the Penrose process can be realized more easily.
@@
Our goal consisted in the derivation and analysis of possible types of scenarios in the simplest case of processes with charged particles thus developing general formalism. This implied consideration of a static metric and electric field. The next step is supposed to be inclusion of rotating, electric charge and a magnetic field altogether into consideration. We hope that the corresponding formalism will be useful for solving realistic problems in astrophysics. However, this needs separate treatment.

### v2

@@ IX.1 Case (i), turning point for particle 2 (TP2)
One can check using (93) and (94) that we can also rewrite $X_{2}$ in the form
$X_{2}=\frac{b_{2}}{2m_{0}^{2}}X_{0}-\frac{\sqrt{d}}{2m_{0}^{2}}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}\text{.}$ (97)
IX.2 Case (i), turning point for particle 0 only (TP0)
@@
Let particles 1 and 2 be ejected in point $r_{d}$ along the trajectory of particle 0. Now, it follows from (48) that either $L_{0}=0$ or $P_{1}=0=P_{2}$. By assumption, the second situation is now impossible and will be considered in the next subsection. Now, we put $L_{0}=0$, whence $\tilde{m}_{0}=m_{0}$, so
@@
IX.3 Case (ii), turning point for all three particles (TP3)
@@
$E_{2}=q_{2}\varphi+\frac{b_{2}}{2m_{0}^{2}}\tilde{m}_{0}N+\frac{\sqrt{d}\left|L_{0}\right|}{2m_{0}^{2}r_{d}}N\text{,}$ (112)
$E_{1}=q_{1}\varphi+\frac{b_{1}}{2m_{0}^{2}}\tilde{m}_{0}N-\frac{\sqrt{d}\left|L_{0}\right|}{2m_{0}^{2}r_{d}}N\text{.}$ (113)
@@
$L_{2}=\frac{L_{0}b_{2}+signL_{0}\sqrt{d}\tilde{m}_{0}r_{d}}{2m_{0}^{2}}=L_{2}^{(+)},$ (114)
$L_{1}=\frac{L_{0}b_{1}-signL_{0}\sqrt{d}\tilde{m}_{0}r_{d}}{2m_{0}^{2}}=L_{1}^{(-)}\text{,}$ (115)
@@
X Threshold for the Penrose process
@@
XI Velocities and gamma factors for scenario I
@@
$\gamma_{01}=\frac{b_{1}}{2m_{0}m_{1}}\text{, }v_{1}=\frac{\sqrt{d}}{b_{1}}\text{,}$ (127)
$V_{1}=\frac{v_{1}+V_{0}}{1+V_{1}V_{0}}\text{,}$ (128)
@@
XII Scenario B and the maximum of efficiency
@@
For particle 1 now
$V_{1}=\frac{v_{1}-V_{0}}{1-V_{1}V_{0}}\text{,}$ (145)
Here, as before, particle 2 moves to infinity (now in the same direction as particle 0) and particle 1 moves in the inward direction (opposite to particle 0). However, now the signs before the radicals are opposite to those in (132), (133). The condition for the PP gives us now
@@
XII.1 Near-horizon limit for scenario B
It is worth noting what happens in the near-horizon limit $N\rightarrow 0$ within the scenario under discussion. On the first glance, it follows from (142) that $X_{2}\equiv E_{2}-q_{2}\varphi$ in this limit can be arbitrary nonzero. However, this is not the case. The point is the correlation between initial conditions and the type of a horizon. Let a particle move away from the nonextremal horizon with finite nonzero $X_{2}$ (so-called usual particle). Then, if we continue its trajectory in the past, it turns out that it appeared there some small proper time $\tau$ ago from the region behind the horizon. But this would be a white, not a black hole and is beyond the scope of our work. See also on details is (there, $q=0$ but this does not matter in the context under discussion).
From the other hand, if the proper time required for crossing the horizon $\tau\rightarrow\infty$, such arguments do not work. This happens if the horizon is extremal. Then, a particle can move in the outward direction from the immediate vicinity of the horizon. But, in doing so, it must have $X=O(N)$ (see frontal for details). Returning to our issue, we see from (142) that both $X_{0}=O(N)$ and $X_{2}=O(N)$.
XIII Efficiency, proximity to horizon and type of scenario
Usually, when considering decay, it is assumed that it happens in the turning point of all particles. In doing so, it is often stated that the efficiency reaches its maximum if decay occurs near the horizon (see, e.g. the review fifty ). Meanwhile, these statements are not quite accurate and require some essential reservations. Also, they apply to scenario TP3 but, in general, not to all other ones. As this concerns the important aspects of process near a black hole, this needs more careful discussion that is given below. Let us consider different types of scenarios in this context case by case.
XIII.1 Scenario I
It is shown above that, according to (60) this scenario requires $X_{0}\leq\frac{N_{d}b_{2}}{2m_{2}}$ in the point of decay $r_{d}.$ This means that for a usual (not fine-tuned) particle that has $X_{0}\neq 0$ on the horizon separated from zero, the horizon limit cannot be taken at all, provided $m_{2}$ is a massive particle. If it has almost vanishing mass, $m_{2}=\mu_{2}N_{d}$ with $\mu_{2}\neq 0$, the situation changes since (60) gives us
$X_{0}\leq\frac{b_{2}}{2\mu_{2}}.$ (153)
If this criterion is fulfilled, the horizon limit is indeed possible for a usual particle. Otherwise, fine-tuning of particle 0 is required.
Let, for simplicity, particle 0 be neutral. Then, if
$q_{2}>q_{2}^{\ast}=\frac{\sqrt{M^{2}-Q^{2}}}{E_{0}Q}\sqrt{d}r_{+}\text{,}$ (154)
the maximum is indeed achieved on the horizon (see XVIII for details). At the same time, condition (117) should be fulfilled as well. On the horizon it reduces to (177).
However, if $q_{2}<q_{2}^{\ast}$, the horizon corresponds not to the maximum but to a local minimum of $E_{2}$.
If $m_{2}=0$ exactly, the horizon limit is also possible but according to (74) this entails a quite strong condition $X_{2}=O(N_{d}^{2})$.
XIII.2 Scenario II
Here, there are no difficulties with decay near the horizon since inequality (70) can be satisfied easily. However, in this scenario both particles fall in a black hole, so this option ceases to be  ”profitable”.
XIII.3 Scenario B
Now, it follows from (142) that for any $q_{2}Q>0$, $\frac{\partial E_{2}}{\partial r}<0$, so maximum is indeed achieved on the horizon. However, another difficulty comes into play here. A usual particle cannot start its motion near the horizon in the outward direction - see XII.1 above, Sec. IA in is and frontal . This is allowed for fine-tuned or near-fine-tuned particles only, with $X_{2}=O(N_{d})$. The same reasonings apply to scenario TP3.
XIII.4 Scenario TP2
According to (94),
$E_{2}=\frac{q_{2}Q}{r}+m_{2}N.$ (155)
Let us, for simplicity, consider the case of the extremal black hole, so $N=1-\frac{r_{+}}{r},Q=M=r_{+}$. Then,
$\frac{\partial E_{2}}{\partial r}=\frac{r_{+}}{r^{2}}(m_{2}-q_{2})\text{.}$ (156)
Thus the biggest value of $E_{2}$ is achieved on the horizon under the condition $m_{2}<q_{2}$ only.
XIII.5 Scenario TP0
According to (105),
$E_{2}=\frac{q_{2}Q}{r}+\frac{b_{2}}{2m_{0}}N\text{.}$ (157)
Then, for the extremal black hole we have
$\frac{\partial E_{2}}{\partial r}=\frac{r_{+}}{r^{2}}(\frac{b_{2}}{2m_{0}}-q_{2})\text{.}$ (158)
The horizon corresponds to the biggest value of $E_{2}$ for $q_{2}>\frac{b_{2}}{2m_{0}}$ only.
We see that one should be very careful making the statement about maximum of efficiency. This necessarily includes indication of scenario and reservation about relation between parameters.
XIV Different scenarios: comparison of properties
Now, it is convenient to summarize the main features of all scenarios in Table I.
Scenario Particle 0 Particle 2 $\delta$ $\left(v_{1}\right)_{\min}$ mandatory $X_{2}$ near horizon
I in out $-1$ no $O(N)$ for $m_{2}\neq 0$, $O(N^{2})$ for $m_{2}=0$
II in in $+1$ yes $O(1)$
TP2 in 0 $-1$ no $O(N)$
TP0 0 out no $O(N)$
TP3 0 0 ±$\pm 1$ no if $\delta=-1$, yes if $\delta=+1$ $O(N)$
B out out $+1$ no $O(N)$
Table I. Classification and main features of scenarios.
We included in it scenarios in which particles 1 and 2 are ejected along the trajectory of particle 0. However, this is done with one exception in a degenerate case. In scenario TP0 not only $P_{0}=0$ but also $L_{0}=0$ see Sec. IX.2. Then, all the components of the three-momentum vanish, so particle 0 is in rest in this point and there is no tangent vector to the trajectory.
We can write a unifying formula for the energy of particle 2. If it moves in the outward direction after decay of particle 0, it can, in principle, escape to infinity and is potentially subject to the PP. Otherwise, it falls in a black hole.
$E_{2}=q_{2}\varphi+X_{0}\frac{b_{2}}{2m_{0}^{2}}+\delta\frac{\sqrt{d}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}}{2m_{0}^{2}}.$ (159)
If immediately after decay particle 2 moves in the outward/inward direction, we use shortening ”out”/”in”. If it is in the radial turning point, we write ”0”. In scenario TP0 the factor $\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}=0$, so $\delta$ is irrelevant.
The full trajectory of particle 2 is model-dependent and cannot be found without specifying the metric. The efficiency $\eta=\frac{E_{2}}{E_{0}}$. In the 4th column we indicate whether or not some low bound $v_{1}\geq\left(v_{1}\right)_{\min}>0$ is required for the PP to occur. In scenario II particle 2 falls in a black hole, so the low bound on $v_{1}$ is pure formal since it has nothing to do with the PP. The general feature consists in that there is such a bound for $\delta=+1$ and it is absent if $\delta=-1$.
In principle, a combined scenario is also possible. If in a point of decay both particles moves in the inward direction (scenario II) but after bouncing back in the turning point particle 2 changes direction and moves outwardly. In particular, this can happen in the near-horizon region of the extremal black hole with $X_{2}=O(N)$ rn , fh .
In all cases, formally $E_{2}\rightarrow\infty$ when $q_{2}\rightarrow\infty$. However, in realistic situations $q_{2}$ is bounded rn , fh .
XV Discrepancy with two previous works on the subject
The Penrose process with charged particles was also considered in ju where an additional assumption that the black hole charge $Q$ is small was made. Our results do not agree. The authors of ju , according to their eqs. (34), (35), obtained that the outgoing particle after decay near the horizon has a finite nonzero energy. Meanwhile, it follows from our formulas that in this case the energy of an escaping particle (if we neglect $Q$) tends to zero. If we take the charge $Q$ into account, in a similar way the difference $X_{2}=E_{2}-q_{2}\frac{Q}{r_{+}}$ tends to zero. Actually, the subtlety in the issue under discussion consists in the necessity to take into account correlation between dynamics and kinematics. This means that in scenario I (where particle 2 escapes) the sign before the radical in (52) comes with minus, not plus. Instead, one can consider scenario II with finite $X_{1}$ and $X_{2}$ where one of particles has the sign plus but this particle falls in a black hole and does not escape. These features can be seen in Table I.
One more attempt of considering the PP for charged particles was made in sh with a magnetic field $B$ taken into account. In our view, the results for efficiency $\eta=\frac{E_{2}}{E_{0}}$ (in our notations) described by Eqs. (31) and (32) of Ref. sh are incorrect. In the neutral case $q=0$, $Q=0$ and $B=0$ they give $\eta=0$ instead of the Schwarzschild value $\eta_{Sch}$. Therefore, for the RN metric, it also does not reproduce the formulas like (110) of the present paper.
The reason of discrepancy can be explained as follows. The results (31), (32) of sh are based on their eq. (30) that contains the angular velocity of a particle. Therefore, account for angular momentum is necessary. For a decay, say, in the turning point of radial motion for all three particles, the correct values are given by our eqs. (114), (115). Meanwhile, in sh all angular momenta are put to be zero. This leads to contradiction.
Alternatively, one may consider a scenario in which all particles move radially with $L_{0}=L_{1}=L_{2}=0$. However, in this case eqs. (19), (20) of sh from which (31) and (32) were derived, loose their sense since (19) is obtained for motion along a circle. Again, we obtain contradiction.
XVI Relation to the BSW effect
@@
XVII Summary and Conclusions
@@
We formulated the results in two forms. The first one includes the dependence of the outcome on masses of particles. The second form expresses them in terms of velocities and gamma factors of relative motion between a parent particle and daughter ones. The corresponding formulas obtained for a static charged metric are similar to whose obtained in wald for rotating metrics and neutral particles. Meanwhile, there are some essential differences. For the Penrose process to occur in the Kerr or other rotating metric, there is a lower bound on velocity of a daughter particle in the center of mass frame, this bound being rather high wald . This creates obstacles for using this process. Meanwhile, in our case, the situation is more diverse and there are scenarios in which the velocity obeys the upper (not lower) bound, so the Penrose process can be realized more easily.
We discussed in detail an important question how the efficiency of the Penrose process depends on a point in which decay occurred. The results depends in general on a scenario and in this sense classification of scenarios developed in our work enables to elucidate this issue in general setting. One may think that this will be useful for analysis of diverse process near astrophysical black holes since we did not restrict ourselves by a separate scenario or special set of date.
@@
Our goal consisted in the present article not in considering some concrete astrophysical problem but in development of general formalism. We carried out the analysis of possible types of scenarios in the simplest case of processes with charged particles thus developing general approach. This implied consideration of a static metric and electric field. The next step is supposed to be inclusion of rotation, electric charge and a magnetic field altogether into consideration.
We hope that the corresponding formalism will be useful for solving realistic problems in astrophysics. This is supposed to include processes in the accretion discs, extraction of energy from magnetized black holes, properties of ionization of neutral particle falling into a black hole, etc. This needs separate treatment. We hope that model-independent approach developed in our previous work eq and the present one, will be useful for a diverse set of such problems.
XVIII Appendix: dependence of efficiency on a point
For simplicity, we consider here a neutral particle 0. Then, $q_{0}=0$, $q_{1}=-q_{2}$ and $X_{0}=E_{0}$. Let after decay the energy $E_{2}$ (and thus efficiency $\eta=E_{2}/E_{0})$ is given by eq. of the type
$E_{2}=q_{2}\frac{Q}{r}+\frac{b_{2}}{2m_{0}^{2}}E_{0}+\frac{\sqrt{d}}{2m_{0}^{2}}\sqrt{E_{0}^{2}-m_{0}^{2}N^{2}}\text{.}$ (172)
Here, $q_{2}>0$. As $\frac{dN}{dr}>0$, it follows that $\frac{dE_{2}}{dr}>0$ and the maximum is formally achieved for decay on the horizon but with all necessary reservations made in Sec. XIII above. More involved case arises if
$E_{2}=q_{2}\frac{Q}{r}+\frac{b_{2}}{2m_{0}^{2}}E_{0}-\frac{\sqrt{d}}{2m_{0}^{2}}\sqrt{E_{0}^{2}-m_{0}^{2}N^{2}}\text{.}$ (173)
Then, direct calculation shows that on the horizon
$\left(\frac{dE_{2}}{dr}\right)_{+}=\frac{1}{r_{+}^{2}}(\frac{\sqrt{M^{2}-Q^{2}}}{E_{0}}\sqrt{d}-q_{2}Q)\text{.}$ (174)
Thus if the biggest value of $E_{2}$ is required to occur on the horizon, it is necessary that $\left(\frac{dE_{2}}{dr}\right)_{+}<0$, whence
$q_{2}>q_{2}^{\ast}=\frac{\sqrt{M^{2}-Q^{2}}}{E_{0}Q}\sqrt{d}.$ (175)
However, if
$q_{2}<q_{2}^{\ast}\text{,}$ (176)
the horizn crresponds to the smallest value.
Simultaneously, the threshold for the PP (117) gives us on the horizon for $\left|q_{1}\right|=q_{2}$
$q_{2}\geq\frac{E_{0}r_{+}}{2m_{0}^{2}Q}(b_{1}+\sqrt{d}),$ (177)
where we have taken into account that now $\left|q_{1}\right|=q_{2}$. Both inequalities (176) and (177) are compatible with each other, provided
$\frac{E_{0}^{2}}{m_{0}^{2}}\leq 2\frac{\sqrt{d}\sqrt{M^{2}-Q^{2}}}{(b_{1}+\sqrt{d})r_{+}}.$ (178)